if the perimeter of a flag is 460 feet and the width is 70 more than the length what is the length of the flag??????????????????????????????????????????????????
Answer by macston(5194)

P=perimeter=2(L+W); L=Length; W=width=L+70 ft
460 ft=2(L+W) Divide each side by 2.
230 ft=L+W Substitute for W.
230 ft=L+L+70 ft
230 ft=2L+70 ft Subtract 70 ft from each side.
160 ft=2L Divide each side by 2.
80 ft=L ANSWER 1 The length is 80 ft.
W=L+70 ft=80 ft+70 ft=150 ft ANSWER 2: The width is 150 ft.
CHECK:
Perimeter=2(L+W)
460 ft=2(80 ft + 150 ft)
460 ft=2(230 ft)
460 ft=460 ft